OpenStudy (anonymous):

A 285g ball falls vertically downward, hitting the floor with a speed of 2.8 m/s and rebounding upward with a speed of 2.4 m/s (a) find the magnitude of the change in the ball's momentum. (b) find the change in the magnitude of the ball's momentum. (c) which of the two quantities calculated in part A and B is more directly related to the net force acting on the ball during its collision with the floor? (d) explain answer for part C.

OpenStudy (anonymous):

a) magnitude of change is final - initial = 2.4 - (-2.8) = 5.2 m/s p = mv = .285 x 5.2 = 1.482 kgm/s b) change in magnitude is 2.4 - 2.8 = -0.4 m/s p = -2.4 x .285 = .684 kgm/s c) a d) impulse is related to force and change of momentum so a) is more relevant. b) does not take into account the change in direction which is significant.